# Spark Transformer: Reactivating Sparsity in FFN and Attention

### v1

@@ Appendix B Implementation Details on Sparse Matrix Multiplications
With batch size of 1, both Spark FFN and Spark Attention utilize two types of sparse vector-matrix multiplication: vector-masked matrix multiplication and sparse vector-matrix multiplication (Figure 4). Given a vector $q$ and a matrix $w$, vector-masked matrix multiplication multiplies $q$ with the non-masked columns of $w$ based on a masking vector $m$. Masked columns yield a zero output. Sparse vector-matrix multiplication, on the other hand, involves a vector that contains many zeroes being multiplied by a dense matrix.
In Spark FFN, we perform vector-masked matrix multiplication for ${\bm{K}}_{2}^{\top}{\bm{q}}[r\!:]$ (Figure 4(a)). The masking vector is generated from the output of $\operatorname{{Statistical-Top}}_{k}({\bm{K}}_{1}^{\top}{\bm{q}}[:\!r])$. Based on the mask, Spark FFN skips loading the masked columns of $w$ from DRAM (in CPU setup) or HBM (in GPU setup) and the associated computations. On CPU, Spark FFN utilizes SIMD operations (as in the original Gemma implementation). To further enhance performance, Spark FFN utilizes software CPU prefetching ($builtin\_prefetch$) to overlap loading from DRAM to the CPU cache with computations. On GPU, Spark FFN utilizes customized CUDA kernel.
The same masking vector also identifies the zero entries in the intermediate vector that is multiplied by matrix $V$ (Figure 4(b)). For this sparse vector-matrix multiplication, we store the matrix in row format. Each CPU thread (or GPU warp) processes a tile of the matrix while skipping the loading and computation of the masked rows. Prefetching and SIMD operations are applied similarly on CPU in this context.
@@ C.1 Distribution of Inputs to Statistical Top-$k$
The underlying assumption for statistical top-$k$ is that the activation vector upon which it is applied to, namely, the pre-GELU activation in Spark FFN and the pre-softmax activation in Spark Attention, can be modeled as being drawn from an i.i.d. Gaussian distribution. Here we provide empirical evaluation on the distribution of these activation vectors for Spark Transformer. Results for Spark FFN and Spark Attention are provided in Figure C.4 and Figure C.5, respectively. The results show that the distribution holds close proximity to a Gaussian, hence justifying the use of statistical top-$k$.
@@ C.2 Sparsity Level During Evaluation
Complementing Figure 3 which reports sparsity level during pretraining, here we report the sparsity level during evaluation to confirm that statistical top-$k$ produces the same level of sparsity during test time. The results are presented in Figure C.1 for some arbitrarily selected tokens. For Attention, in particular, we select tokens at the positions 512, 1024, and 2048 which are all above our choice of $k=256$ for Spark Attention.
@@
Figure C.1: Sparsity in the intermediate activation of Spark FFN and Spark Attention during evaluation (see Figure 3 for results during training). For FFN, we use a simple prompt “test” and report the percentage of nonzero entries in generating the 5th, 10th, and 15th token. For Attention, we report the nuber of nonzero entries at the 512th, 1024th, and 2048th token during prefill.
@@ C.3 Batching Analysis
Figure C.2 provides the performance comparison between Spark Gemma 2 and Gemma 2, measured in prefill throughput (tokens/sec) across varying chunk sizes. We use a 4096-token prompt on a 16-core CPU VM. A similar trend is expected during the decoding phase with varying batch sizes.
Our analysis shows that Spark Transformer provides the highest speedup at batch size 1, and again at large batch sizes (e.g. $>$8), where the compute FLOP becomes the primary bottleneck.
@@
Spark FFN comes with two hyper-parameters, namely $r$ which controls the rank hence FLOP count of the low-cost predictor, and $k$ which controls sparsity of activation hence the FLOP count. In Figure C.3 we provide an ablation study on the effect of these two hyper-parameters, by reporting the training loss curves in the first 25,000 training steps (which is around 5% of full training). From Figure 3(a), the best choice of $r$ is 1024 which is nearly half of $d_{\text{model}}=2304$ (due to model sharding constraint, $r$ cannot be taken to be exactly a half of $d_{\text{model}}$). From Figure 3(b), we see that the model quality is insensitive to choices of $k$ that gives $[5\%,10\%]$ sparsity, but there is quality loss if we go sparser, e.g. 3% nonzeros.
@@
Figure C.4: Distribution of the entries of the input activation to statistical top-$k$ in Spark FFN (see Figure C.5 for result of Spark Attention). The two rows correspond to activation at two positions 0 and 1000 of an input, and the columns correspond to activation at four different depth levels $\{0,8,16,24\}$ of the 26-layer pretrained Spark Transformer. The input is the first 1000 tokens of the first essay from https://paulgraham.com/articles.html prepended with the BOS token. We compare the empirical distribution (Empirical) with the Gaussian distribution whose mean and standard deviation (std) are computed as the sample mean and std of the input (Fitted). We see that the Gaussian closely approximates the empirical distribution. We also compare the cutoff value estimated from the Gaussian, i.e., $\theta({\bm{x}},k)$ used in Equation 10 with $k/d=5\%$ (Cutoff for fitted), with the cutoff value for obtaining $8\%$ nonzeros on the empirical distribution (Cutoff for empirical). It can be seen that these two cutoff values are close.
@@
Figure C.5: Distribution of the entries of the input activation to statistical top-$k$ in Spark Attention (see Figure C.4 for result of Spark FFN). The two rows correspond to activation for two different attention heads, and the columns correspond to activation at four different depth levels $\{0,8,16,24\}$ of the 26-layer pretrained Spark Transformer. Model input is the first 1000 tokens of the first essay from https://paulgraham.com/articles.html prepended with the BOS token, and we examine activation of the last token (i.e., inner product between the query embedding of the 1001st token and all 1001 key embeddings). We compare the empirical distribution (Empirical) with the Gaussian distribution whose mean and standard deviation (std) are computed as the sample mean and std of the input (Fitted). We see that the Gaussian closely approximates the empirical distribution. We also compare the cutoff value estimated from the Gaussian, i.e., $\theta({\bm{x}},k)$ used in Equation 10 with $k=256$ (Cutoff for fitted), with the cutoff value for obtaining top 256 entries on the empirical distribution (Cutoff for empirical). It can be seen that these two cutoff values are close.
@@ C.5 Additional Ablation Studies
In this section, we provide ablation studies for understanding the effect of the individual components of Spark Transformer. Towards that, we plot the training loss curves for Gemma-2 and Spark Transformer, see Figure C.6. Here, we restrict to the first 80k training steps out of the 500k total steps since it is costly to fully train all ablation models, and that 80k steps is sufficient for seeing the trend. We can see that Spark Transformer slightly lags behind Gemma-2. However, as demonstrated in Figure 1(b), that small difference in training loss does not lead to a substantial difference in evaluation quality.
@@
To understand the effect of Spark FFN vs Spark Attention, we conduct an experiment where only attention is switched from a standard one to Spark Attention, whereas the FFN remains the standard one. The result is illustrated as Gemma-2 + Spark Attention in Figure C.6. It can be seen that Spark Attention provides a minor quality gain over Gemma-2. In comparing Gemma-2 + Spark Attention with Spark Transformer, this also shows that further adding Spark FFN slightly hurts model quality. As noted above, such a small difference does not lead to substantial quality impact on the evaluation tasks. Hence, we conclude here that none of Spark FFN and Spark Attention has significant quality impact.
@@
Sparsity enforcing via statistical top-$k$ and low-cost activation predictor are two relatively independent components of Spark Transformer. This means that, upon the standard Gated FFN (see Equation 5) that is used in Gemma-2, which we rewrite here for convenience:
@@
Note that applying a sparsifying function on the input to the nonlinear function $\sigma()$ as in Equation C.2 is a common choice in the literature of sparse activations, e.g., [59, 76, 45]; the main difference between these works lies in the specific sparsity enforcing technique, see Table D.1 for a summary. In addition to the sparsifying function, Spark FFN also has another architectural change for the purpose of introducing a low-cost predictor. Here, we rewrite Spark FFN for ease of comparison with Equation C.2:
@@
Analogous to FFN, we may also only add statistical top-$k$ to attention without the low-cost predictor, i.e., by switching from standard Attention in Equation 6 to the following:
@@
Here, we aim to understand the effect of introducing statistical top-$k$ without the low-cost predictor. Towards that, we conduct an experiment where FFN and Attention in Gemma-2 are replaced with Equation C.2 and Equation C.4, respectively. The result is illustrated as Gemma-2 + Top-k in Figure C.6. It can be seen that the training loss becomes notably larger and the gap compared to Gemma-2 is further increasing with more training steps. This result demonstrates that while the low-cost predictor is introduced to Spark Transformer for reducing the cost in predicting the nonzero entries, it also helps in bridging the gap from the introduction of statistical top-$k$. In other words, Transformer with low-rank predictors in FFN and Attention is more amenable to activation sparsification without quality loss.
@@ C.6 Computing FLOPs per Token
To understand the FLOPs saving reported in Figure 1(a), we provide a comparison between FLOPs of the major components of a Transformer, including FFN, Attention dot product, and Attention projections, with that of Spark Transformer. The results are presented in Table C.1.
Table C.1: FLOPs per token comparison: Spark Transformer vs. standard Transformer. In a standard Transformer with model dimension $d_{\text{model}}$, we assume multi-head attention where the sum of head dimensions equals $d_{\textnormal{model}}$, and an FFN with non-gated activation and width $d_{\textnormal{ff}}$. Here, $n_{\textnormal{ctx}}$ represents the context length for the target token. The computational cost is primarily determined by the FFN (assuming $d_{\textnormal{ff}}\gg d_{\textnormal{model}}$, which is typical) and the attention dot product (assuming a long context length). Spark Transformers introduce sparsity parameters, $k_{\textnormal{ff}}$ and $k_{\textnormal{attn}}$, to reduce FLOPs. Setting $k_{\textnormal{ff}}=8\%\times d_{\textnormal{ff}}$ and $k_{\textnormal{attn}}=256$ achieves a 3.2$\times$ FLOPs reduction in the FFN, a 4$\times$ reduction in the attention dot product, and a 2.5$\times$ reduction overall (assuming $n_{\textnormal{ctx}}=8k$) for Gemma-2B.
@@ D.1 Comparison with Related Work on Activation Sparsity in FFN
In Table D.1, we provide a summary of recent work on enabling FFN activation sparsity in the latest LLMs, including ReLUification [59], ProSpare [76], HiRE [83], and CATS [45]. Please see Section 5 for a discussion of these methods.
@@
  • Pretraining-based approaches are those whose activation sparsity is enforced from the very beginning of the model pretraining, usually with certain architectural changes upon an established model. Hence, such methods require pretraining a model from scratch. For example, ReLUification [59] is the method of switching the activation function in FFN from a commonly used one, e.g. GELU, back to ReLU. When tested on OPT 1.3B, [59] applied this modification and pretrained a version of OPT 1.3B with ReLU from scratch (Table D.1 contains a summary of relevant results for it).
  • Finetuning-based approaches refer to those that take an existing pretrained model without activation sparsity, and perform additional training steps usually with certain architectural changes to induce sparsity. For example, ReLUification discussed above as a pretraining-based approach can also be used as a finetuning-based approach by switching the activation function of a pretrained model to ReLU and performing additional training steps. In [59], this approach is applied to Falcon 7B and Llama 7B, for which the results are summarized in Table D.1. Other methods falling into this category, which we have summarized in Table D.1, include ProSparse [76] and HiRE [83].
  • “Zero-shot” approaches. This refers to methods that enforces activation sparsity upon a pretrained model, without any additional training. CATS [45] is a method of this kind.
Our Spark Transformer belongs to the category of pretraining-based method, that is, it requires pre-training a model from scratch. It may not be used directly as a finetuning-based approach, due to the drastic difference between the architecture of a standard FFN with gated activation, i.e. Equation 5, and our Spark-FFN. This appears to make Spark Transformer less favorable than finetuning-based approaches, as the latter requires usually a small fraction of the pretraining steps in the finetuning process, see Table D.1 for a summary. However, the long-term serving costs of such models usually dominate overall expenditures, in which case the additional pre-training costs are amortized by the significant efficiency benefits realized from a higher level of sparsity and FLOPs reduction during model inference.
@@
ReLUification [^7] Results reported here are for the stage 1 training of their paper. [59] ReLU None OPT 1.3B +0% 93% -2% -62%
@@
ProSparse [76] ReLU + $\|\cdot\|_{1}$ None Llama2 7B
@@ D.2 Additional Discussions on Statistical Top-$k$
Novelty upon existing work. We note that ideas similar to our statistical top-$k$ have appeared in the literature. In particular, [74] introduced the idea of fitting a Gaussian distribution to the entries of an input vector and estimating a threshold from quantile functions. Then, [57] extended the approach to additional distributions. In both cases, the method is used for solving the problem of distributed training. Here, we summarize our contribution upon these works:
  • We are the first to use statistical top-$k$ for enforcing activation sparsity in Transformers. Improving Transformer efficiency via activation sparsity has become a very popular research topic (see Section 5), but may have been suffering from a lack of efficient top-$k$ algorithms for enforcing sparsity. Hence, the introduction of statistical top-$k$ may facilitate the development of this area.
@@
  • We reveal the conceptual connection between statistical top-$k$ and several related top-$k$ operators in the literature, see Section 3.3. Such connections may motivate the development of more powerful top-k algorithms in the future.
@@
  • Global statistical top-$k$. Here we require each device to compute the mean and variance for entries on itself, then communicate them to all other devices. In this case, each device receives $m-1$ mean and variance values, which can be combined with mean and variance on its own to obtain global mean and global variance. Then, the rest of the steps in statistical top-$k$ can be carried out on individual devices. In this method, the output is exactly the same as if applying statistical top-$k$ without activation sharding. In terms of cost, there is extra computation coming from aggregating mean and variance from all devices, but the cost is very low as it requires only $O(k)$ FLOPs. The method also introduces a communication cost, but the cost is again small as each device only needs to send / receive $22k-$ 2 floating point numbers.

### v2

@@ Appendix B Implementation Details on Sparse Matrix Multiplications
With batch size of 1, both Spark FFN and Spark Attention utilize two types of sparse vector-matrix multiplication: vector-masked matrix multiplication and sparse vector-matrix multiplication (Figure˜4). Given a vector $q$ and a matrix $w$, vector-masked matrix multiplication multiplies $q$ with the non-masked columns of $w$ based on a masking vector $m$. Masked columns yield a zero output. Sparse vector-matrix multiplication, on the other hand, involves a vector that contains many zeroes being multiplied by a dense matrix.
In Spark FFN, we perform vector-masked matrix multiplication for ${\bm{K}}_{2}^{\top}{\bm{q}}[r\!:]$ (Figure˜4(a)). The masking vector is generated from the output of $\operatorname{{Statistical-Top}}_{k}({\bm{K}}_{1}^{\top}{\bm{q}}[:\!r])$. Based on the mask, Spark FFN skips loading the masked columns of $w$ from DRAM (in CPU setup) or HBM (in GPU setup) and the associated computations. On CPU, Spark FFN utilizes SIMD operations (as in the original Gemma implementation). To further enhance performance, Spark FFN utilizes software CPU prefetching (builtin_prefetchbuiltin\_prefetch) to overlap loading from DRAM to the CPU cache with computations. On GPU, Spark FFN utilizes customized CUDA kernel.
The same masking vector also identifies the zero entries in the intermediate vector that is multiplied by matrix $V$ (Figure˜4(b)). For this sparse vector-matrix multiplication, we store the matrix in row format. Each CPU thread (or GPU warp) processes a tile of the matrix while skipping the loading and computation of the masked rows. Prefetching and SIMD operations are applied similarly on CPU in this context.
@@ C.1 Distribution of Inputs to Statistical Top-$k$
The underlying assumption for statistical top-$k$ is that the activation vector upon which it is applied to, namely, the pre-GELU activation in Spark FFN and the pre-softmax activation in Spark Attention, can be modeled as being drawn from an i.i.d. Gaussian distribution. Here we provide empirical evaluation on the distribution of these activation vectors for Spark Transformer. Results for Spark FFN and Spark Attention are provided in Figure˜C.4 and Figure˜C.5, respectively. The results show that the distribution holds close proximity to a Gaussian, hence justifying the use of statistical top-$k$.
@@ C.2 Sparsity Level During Evaluation
Complementing Figure˜3 which reports sparsity level during pretraining, here we report the sparsity level during evaluation to confirm that statistical top-$k$ produces the same level of sparsity during test time. The results are presented in Figure˜C.1 for some arbitrarily selected tokens. For Attention, in particular, we select tokens at the positions 512, 1024, and 2048 which are all above our choice of $k=256$ for Spark Attention.
@@
Figure C.1: Sparsity in the intermediate activation of Spark FFN and Spark Attention during evaluation (see Figure˜3 for results during training). For FFN, we use a simple prompt “test” and report the percentage of nonzero entries in generating the 5th, 10th, and 15th token. For Attention, we report the nuber of nonzero entries at the 512th, 1024th, and 2048th token during prefill.
@@ C.3 Batching Analysis
Figure˜C.2 provides the performance comparison between Spark Gemma 2 and Gemma 2, measured in prefill throughput (tokens/sec) across varying chunk sizes. We use a 4096-token prompt on a 16-core CPU VM. A similar trend is expected during the decoding phase with varying batch sizes.
Our analysis shows that Spark Transformer provides the highest speedup at batch size 1, and again at large batch sizes (e.g. >>8), where the compute FLOP becomes the primary bottleneck.
@@
Spark FFN comes with two hyper-parameters, namely $r$ which controls the rank hence FLOP count of the low-cost predictor, and $k$ which controls sparsity of activation hence the FLOP count. In Figure˜C.3 we provide an ablation study on the effect of these two hyper-parameters, by reporting the training loss curves in the first 25,000 training steps (which is around 5% of full training). From Figure˜3(a), the best choice of $r$ is 1024 which is nearly half of $d_{\text{model}}=2304$ (due to model sharding constraint, $r$ cannot be taken to be exactly a half of $d_{\text{model}}$). From Figure˜3(b), we see that the model quality is insensitive to choices of $k$ that gives $5\%,10\%]$ sparsity, but there is quality loss if we go sparser, e.g. 3% nonzeros.
@@
Figure C.4: Distribution of the entries of the input activation to statistical top-$k$ in Spark FFN (see Figure˜C.5 for result of Spark Attention). The two rows correspond to activation at two positions 0 and 1000 of an input, and the columns correspond to activation at four different depth levels $\{0,8,16,24\}$ of the 26-layer pretrained Spark Transformer. The input is the first 1000 tokens of the first essay from https://paulgraham.com/articles.html prepended with the BOS token. We compare the empirical distribution (Empirical) with the Gaussian distribution whose mean and standard deviation (std) are computed as the sample mean and std of the input (Fitted). We see that the Gaussian closely approximates the empirical distribution. We also compare the cutoff value estimated from the Gaussian, i.e., $\theta({\bm{x}},k)$ used in Equation˜10 with $k/d=5\%$ (Cutoff for fitted), with the cutoff value for obtaining $8\%$ nonzeros on the empirical distribution (Cutoff for empirical). It can be seen that these two cutoff values are close.
@@
Figure C.5: Distribution of the entries of the input activation to statistical top-$k$ in Spark Attention (see Figure˜C.4 for result of Spark FFN). The two rows correspond to activation for two different attention heads, and the columns correspond to activation at four different depth levels $\{0,8,16,24\}$ of the 26-layer pretrained Spark Transformer. Model input is the first 1000 tokens of the first essay from %****␣main_arxiv_v3_gdm.tex␣Line␣1475␣****https://paulgraham.com/articles.html prepended with the BOS token, and we examine activation of the last token (i.e., inner product between the query embedding of the 1001st token and all 1001 key embeddings). We compare the empirical distribution (Empirical) with the Gaussian distribution whose mean and standard deviation (std) are computed as the sample mean and std of the input (Fitted). We see that the Gaussian closely approximates the empirical distribution. We also compare the cutoff value estimated from the Gaussian, i.e., $\theta({\bm{x}},k)$ used in Equation˜10 with $k=256$ (Cutoff for fitted), with the cutoff value for obtaining top 256 entries on the empirical distribution (Cutoff for empirical). It can be seen that these two cutoff values are close.
@@ C.5 Additional Ablation Studies
In this section, we provide ablation studies for understanding the effect of the individual components of Spark Transformer. Towards that, we plot the training loss curves for Gemma-2 and Spark Transformer, see Figure˜C.6. Here, we restrict to the first 80k training steps out of the 500k total steps since it is costly to fully train all ablation models, and that 80k steps is sufficient for seeing the trend. We can see that Spark Transformer slightly lags behind Gemma-2. However, as demonstrated in Figure˜1(b), that small difference in training loss does not lead to a substantial difference in evaluation quality.
@@
To understand the effect of Spark FFN vs Spark Attention, we conduct an experiment where only attention is switched from a standard one to Spark Attention, whereas the FFN remains the standard one. The result is illustrated as Gemma-2 + Spark Attention in Figure˜C.6. It can be seen that Spark Attention provides a minor quality gain over Gemma-2. In comparing Gemma-2 + Spark Attention with Spark Transformer, this also shows that further adding Spark FFN slightly hurts model quality. As noted above, such a small difference does not lead to substantial quality impact on the evaluation tasks. Hence, we conclude here that none of Spark FFN and Spark Attention has significant quality impact.
@@
Sparsity enforcing via statistical top-$k$ and low-cost activation predictor are two relatively independent components of Spark Transformer. This means that, upon the standard Gated FFN (see Equation˜5) that is used in Gemma-2, which we rewrite here for convenience:
@@
Note that applying a sparsifying function on the input to the nonlinear function $\sigma()$ as in Equation˜C.2 is a common choice in the literature of sparse activations, e.g., Mirzadeh et al. (2023); Song et al. (2024a); Lee et al. (2024b); the main difference between these works lies in the specific sparsity enforcing technique, see Table˜D.1 for a summary. In addition to the sparsifying function, Spark FFN also has another architectural change for the purpose of introducing a low-cost predictor. Here, we rewrite Spark FFN for ease of comparison with Equation˜C.2:
@@
Analogous to FFN, we may also only add statistical top-$k$ to attention without the low-cost predictor, i.e., by switching from standard Attention in Equation˜6 to the following:
@@
Here, we aim to understand the effect of introducing statistical top-$k$ without the low-cost predictor. Towards that, we conduct an experiment where FFN and Attention in Gemma-2 are replaced with Equation˜C.2 and Equation˜C.4, respectively. The result is illustrated as Gemma-2 + Top-k in Figure˜C.6. It can be seen that the training loss becomes notably larger and the gap compared to Gemma-2 is further increasing with more training steps. This result demonstrates that while the low-cost predictor is introduced to Spark Transformer for reducing the cost in predicting the nonzero entries, it also helps in bridging the gap from the introduction of statistical top-$k$. In other words, Transformer with low-rank predictors in FFN and Attention is more amenable to activation sparsification without quality loss.
@@ C.6 Computing FLOPs per Token
To understand the FLOPs saving reported in Figure˜1(a), we provide a comparison between FLOPs of the major components of a Transformer, including FFN, Attention dot product, and Attention projections, with that of Spark Transformer. The results are presented in Table˜C.1.
[…]
@@ D.1 Comparison with Related Work on Activation Sparsity in FFN
In Table˜D.1, we provide a summary of recent work on enabling FFN activation sparsity in the latest LLMs, including ReLUification (Mirzadeh et al., 2023), ProSpare (Song et al., 2024a), HiRE (Yerram et al., 2024), and CATS (Lee et al., 2024b). Please see Section˜5 for a discussion of these methods.
@@
  • Pretraining-based approaches are those whose activation sparsity is enforced from the very beginning of the model pretraining, usually with certain architectural changes upon an established model. Hence, such methods require pretraining a model from scratch. For example, ReLUification (Mirzadeh et al., 2023) is the method of switching the activation function in FFN from a commonly used one, e.g. GELU, back to ReLU. When tested on OPT 1.3B, Mirzadeh et al. (2023) applied this modification and pretrained a version of OPT 1.3B with ReLU from scratch (Table˜D.1 contains a summary of relevant results for it).
  • Finetuning-based approaches refer to those that take an existing pretrained model without activation sparsity, and perform additional training steps usually with certain architectural changes to induce sparsity. For example, ReLUification discussed above as a pretraining-based approach can also be used as a finetuning-based approach by switching the activation function of a pretrained model to ReLU and performing additional training steps. In Mirzadeh et al. (2023), this approach is applied to Falcon 7B and Llama 7B, for which the results are summarized in Table˜D.1. Other methods falling into this category, which we have summarized in Table˜D.1, include ProSparse (Song et al., 2024a) and HiRE (Yerram et al., 2024).
  • “Zero-shot” approaches. This refers to methods that enforces activation sparsity upon a pretrained model, without any additional training. CATS (Lee et al., 2024b) is a method of this kind.
Our Spark Transformer belongs to the category of pretraining-based method, that is, it requires pre-training a model from scratch. It may not be used directly as a finetuning-based approach, due to the drastic difference between the architecture of a standard FFN with gated activation, i.e. Equation˜5, and our Spark-FFN. This appears to make Spark Transformer less favorable than finetuning-based approaches, as the latter requires usually a small fraction of the pretraining steps in the finetuning process, see Table˜D.1 for a summary. However, the long-term serving costs of such models usually dominate overall expenditures, in which case the additional pre-training costs are amortized by the significant efficiency benefits realized from a higher level of sparsity and FLOPs reduction during model inference.
@@
ReLUification [^8] Results reported here are for the stage 1 training of their paper. (Mirzadeh et al., 2023) ReLU None OPT 1.3B +0% 93% -2% -62%
@@
ProSparse (Song et al., 2024a) ReLU + $\|\cdot\|_{1}$ None Llama2 7B
@@ D.2 Additional Discussions on Statistical Top-$k$
Novelty upon existing work. We note that ideas similar to our statistical top-$k$ have appeared in the literature. In particular, Shi et al. (2019) introduced the idea of fitting a Gaussian distribution to the entries of an input vector and estimating a threshold from quantile functions. Then, M Abdelmoniem et al. (2021) extended the approach to additional distributions. In both cases, the method is used for solving the problem of distributed training. Here, we summarize our contribution upon these works:
  • We are the first to use statistical top-$k$ for enforcing activation sparsity in Transformers. Improving Transformer efficiency via activation sparsity has become a very popular research topic (see Section˜5), but may have been suffering from a lack of efficient top-$k$ algorithms for enforcing sparsity. Hence, the introduction of statistical top-$k$ may facilitate the development of this area.
@@
  • We reveal the conceptual connection between statistical top-$k$ and several related top-$k$ operators in the literature, see Section˜3.3. Such connections may motivate the development of more powerful top-k algorithms in the future.
@@
  • Global statistical top-$k$. Here we require each device to compute the mean and variance for entries on itself, then communicate them to all other devices. In this case, each device receives m1m-1 mean and variance values, which can be combined with mean and variance on its own to obtain global mean and global variance. Then, the rest of the steps in statistical top-$k$ can be carried out on individual devices. In this method, the output is exactly the same as if applying statistical top-$k$ without activation sharding. In terms of cost, there is extra computation coming from aggregating mean and variance from all devices, but the cost is very low as it requires only $O(k)$ FLOPs. The method also introduces a communication cost, but the cost is again small as each device only needs to send / receive 2k22k-2 floating point numbers.
@@
D.3 Synergy with Speculative Decoding
Speculative decoding is a prominent technique for accelerating the inference of large autoregressive models (Leviathan et al., 2023; Chen et al., 2023). It employs a smaller, faster “draft model” to generate a block of candidate tokens, which are then verified in a single, parallel forward pass by the larger, more accurate “target model”. We view Spark Transformer as a highly complementary and synergistic technology, as its architecture enhances efficiency in both roles within the speculative decoding framework.
Spark Transformer as the target model. The primary computational bottleneck in speculative decoding is the parallel verification step, where the target model processes multiple tokens simultaneously. Spark Transformer is ideally suited to accelerate this step.
As the target model, Spark Transformer directly reduces the latency of this bottleneck, leveraging the same wall-time speedups demonstrated in single-token decoding. While verifying multiple tokens in parallel (e.g., $k=2$ to $k=4$) necessarily activates more neurons than a single-token pass, significant sparsity is still maintained. This is for two primary reasons:
  • Small Draft Size: The number of speculative tokens is typically small, meaning the union of activated neurons remains a small fraction of the total.
  • Aggregated Sparsity: Consecutive tokens in a natural language sequence often share a substantial portion of their activated neurons. This phenomenon, termed aggregated sparsity (Mirzadeh et al., 2023), is an extension of the lazy neuron phenomenon (Li et al., 2022) and the observed persistence of activation patterns for related inputs.
Because the union of activated FFN neurons and attended tokens remains sparse, Spark Transformer effectively reduces the computational cost of the most expensive part of the speculative decoding loop.
Spark Transformer as the Draft Model. An effective draft model must satisfy two criteria: it must be significantly faster than the target model, and it must be accurate (i.e., its predictions must frequently align with the target model’s). High accuracy is crucial as it dictates the token acceptance rate, which is the primary driver of the overall speedup.
Spark Transformer is a strong candidate for a high-performance draft model. Our results demonstrate that it achieves near-quality neutrality with its dense counterpart while operating at a fraction of the inference cost. Using a Spark Transformer as the draft model — either a smaller version or even the same model with a lower sparsity budget — could lead to a much higher acceptance rate than a standard distilled model of equivalent speed. This high fidelity, combined with low latency, makes it an ideal driver for maximizing the efficiency of the speculative decoding process.
D.4 Synergy with Quantization Techniques
We view Spark Transformer and quantization as two orthogonal and highly synergistic optimization axes. Quantization, such as conversion to FP8 or INT8, reduces the computational cost and memory size of each individual operation and parameter. Spark Transformer, in contrast, reduces the total number of operations and memory accesses by dynamically pruning FFN activations and attention tokens. Critically, our sparse implementation also reduces the memory bandwidth bottleneck by skipping the load of masked weights from HBM. The combined benefits are therefore expected to be multiplicative, leading to significant compounded gains in inference throughput and energy efficiency.
A key question is whether a model that is already sparsely activated is more sensitive to the precision loss from quantization. We hypothesize that, for activation quantization in particular, Spark Transformer may reduce the model’s sensitivity and be less prone to accuracy degradation compared to a dense model. This is critical, as recent work has shown that standard sparsity and quantization can be in contradiction: magnitude-based sparsity preserves large-value outliers that are detrimental to quantization-aware scaling (Guo et al., 2024). This hypothesis stems directly from the design of our statistical top-$k$ operator.
  • Dynamic Range Compression: The statistical top-$k$ operator in Spark Transformer employs soft-thresholding. This function outputs $\max(x-\theta(x,k),0)$, effectively subtracting the learned threshold $\theta(x,k)$ from all activated neurons. This operation inherently compresses the dynamic range of the activation tensor. A primary challenge in LLM activation quantization is the presence of large-magnitude "outlier" values, which skew the quantization scaling factor and lead to significant precision loss for the more common, smaller values (Dettmers et al., 2022). By “pre-conditioning” the distribution and shrinking these outliers, our operator makes the activation tensor significantly more amenable to accurate low-bit representation.
  • Zero-Point Stability: The 92% of FFN neurons that are not in the top-k are set to 0. This large volume of true zeros is perfectly and losslessly represented in any quantization format, avoiding the "near-zero" noise that can plague dense models.
For weight quantization, we hypothesize that the sensitivity would be broadly similar to the dense Gemma-2 counterpart, as Spark Transformer reallocates and reuses the full set of parameters rather than pruning them.
While a full empirical study of this interaction is a highly promising direction for future work, we hypothesize that Spark Transformer is not only compatible with quantization but may actively facilitate it. This positions our work alongside other recent efforts to create novel sparse-quantized representations (Dettmers et al., 2023), offering a robust path to further acceleration."